OpenStudy (anonymous):

how would i find the domain and vertical asymptotes for f(x)=x-1/x^2-2x+1?

OpenStudy (anonymous):

set the denominator = 0 and solve , then eliminate the x's from the reals

OpenStudy (anonymous):

The domain is all real number except \[x \neq1\] and the vertical asymptote of the function is at x=1 itself.

OpenStudy (anonymous):

so the vertical is the top while the domain is x-1=x^2-2x+1 and then solve?

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

pls explain

OpenStudy (anonymous):

you should factorize the denominator as the denominator will influence your domain and vertical asymptote.. what mathg8 try saying is that you should put x^2-2x+1 = 0

OpenStudy (anonymous):

oh, then what role does x-1 play?

OpenStudy (anonymous):

vertical i assume?

OpenStudy (anonymous):

You actually need that if you want to find the horizontal asymptote of your function.. but as for your question right there, the (x-1) is kinda useless...

OpenStudy (anonymous):

there are horizontal and slant asymptotes that depend on the numerator

OpenStudy (anonymous):

oh, don't need those then... but what does x^2-2x+1 tell me then...

OpenStudy (anonymous):

here's an easy example : what is the domain of 1/x ?

OpenStudy (anonymous):

all real numbers?

OpenStudy (anonymous):

well ... no ...what happens if you divide by 0?

OpenStudy (anonymous):

besides 0

OpenStudy (anonymous):

yes ! you can't divide by ZERO , that's why we exclude the values of x for which the denominator becomes ZERO

OpenStudy (anonymous):

making the denominator =0 we find the values of x that make the denominator zero ...so we have to eliminate them from the domain

OpenStudy (anonymous):

got it got it

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

thanks for meds...

OpenStudy (anonymous):

my pleasure